OpenStudy (anonymous):

find the range of f(x)= (x^2 + 1)/(x^2 + 5) if the domain is (-1/2<= x<= 2)

OpenStudy (tkhunny):

Oh, that fellow is VERY positive. Can you even get it down to zero?

OpenStudy (aum):

Find f'(x) and the critical point(s) in the domain [-1/2, 2]. Find where the function is increasing/decreasing.

OpenStudy (aum):

\[ f(x)= \frac{(x^2 + 1)}{(x^2 + 5)} \\ f'(x) = \frac{(x^2+5)(2x)-(x^2+1)(2x)}{(x^2+5)^2} = \frac{2x(x^2+5-x^2-1)}{(x^2+5)^2} = \frac{8x}{(x^2+5)^2} = 0\\ x = 0 \text{ is a critical point.} \\ x \lt 0, f'(x) < 0, f(x) \text{ is decreasing.} \\ x \gt 0, f'(x) > 0, f(x) \text{ is increasing.} \\ \]

OpenStudy (aum):

f(x) attains minimum when x = 0 and f(0) = (0+1)/(0+5) = 1/5 = 0.2 In [-1/2, 0), f(x) is decreasing. So maximum at x = -1/2. f(-1/2) = (1/4+1)/(1/4+5) = 5/21. In (0, 2], f(x) is increasing. So maximum at x = 2. f(2) = (4+1)/(4+5) = 5/9. 5/9 > 5/21 So the range of f(x) in the domain [-1/2, 2] is [1/5, 5/9].

OpenStudy (aum):

The question is still open so I assume you have probably not covered derivatives, critical points, etc. yet. You can also find the range algebraically: \[ f(x)= \frac{(x^2 + 1)}{(x^2 + 5)} = \frac{(x^2 + 5-4)}{(x^2 + 5)} = \frac{(x^2 + 5)}{(x^2 + 5)} - \frac{(4)}{(x^2 + 5)} = 1 - \frac{4}{(x^2 + 5)} \\ \]The minimum value of \((x^2+5)\) is when x = 0 and it is 5. So the maximum value of \(\large \frac{4}{(x^2+5)}\) is 4/5 and the minimum value of f(x) is 1 - 4/5 = 1/5. Like before, f(-1/2) = 5/21; f(2) = 5/9. 5/9 > 5/21. So the range of f(x) in the domain [-1/2, 2] is [1/5, 5/9].